OpenStudy (anonymous):

Use basic identities to simplify the expression. cos θ - cos θ sin2θ

OpenStudy (anonymous):

\[\sin(2\theta)=2\sin(\theta)\cos(\theta)\] \[\cos( θ) - cos( θ) sin(2θ)=cos( θ)(1-sin(2θ))\] \[\cos( θ)(1-2\sin(\theta)\cos(\theta))\] \[=cos( θ)-2\sin(\theta)\cos^2(\theta)\]